### v1

# SN 2024abvb: A Type Icn Supernova in the Outskirts of its Host Galaxy

Abstract
We present multiband photometric and spectroscopic observations of supernova (SN) 2024abvb, which exhibits early-time prominent photoionized narrow emission lines of C II superposed on a blue continuum. The absence of Balmer features indicates that the SN exploded within hydrogen-poor circumstellar matter (CSM). Together with the lack of explicit evidence of helium signatures, we tentatively identify SN 2024abvb as a Type Icn SN (SN Icn). After correcting for extinction, we estimate an -band peak absolute magnitude of , placing SN 2024abvb in the luminous regime of SNe Icn. We adopted a hybrid model that accounts for both the energy released by the ejecta-CSM interaction and the radioactive decay of nickel synthesized in the SN ejecta to fit the light curve of SN 2024abvb. The best-fit model to the multiband light curves within the first days after explosion suggests that the CSM, radioactive nickel, and ejecta masses to be , , and , respectively. Such a low ejecta mass indicates that the progenitor star of SN 2024abvb experienced a significant mass-stripping process, consistent with the hydrogen-poor and helium-poor spectral features. SN 2024abvb provides important insights into the physical origins of the rare subclass of SNe Icn.
show]wang_xf@mail.tsinghua.edu.cn
I Introduction
Interacting supernovae (SNe) are a unique type of transient characterized by strong interaction between the SN ejecta and circumstellar matter (CSM; 2017hsn..book..403S). The presence of the latter may obscure the prompt emission from the former to some extent, depending on the details of the ejecta kinematics and the CSM distribution. The resulting spectral signatures manifest as a series of narrow photoionized emission lines superposed on a blue continuum (e.g., 2001MNRAS.325..907F; 2004MNRAS.352.1213C; 2010ApJ...709..856S; 2012AJ....144..131Z; 2014Natur.509..471G; 2016ApJ...818....3K; 2017NatPh..13..510Y; 2017A&A...605A..83D; 2021ApJ...912...46B; 2023ApJ...952..119B; 2024ApJ...970..189J). Substantial efforts have gone into establishing the relationship between the physical properties of the surrounding CSM and the mass-loss history of progenitor stars, challenging the existing paradigms of stellar evolution.
One major hydrogen-deficient type of interacting events involves Type Ibn SNe (SNe Ibn), with a fraction of about 8.8% among the Type Ibc sample(2025A&A...698A.305M). Such transients are characterized by narrow helium emission lines in their spectra, indicating the presence of He-rich and H-poor CSM with an expansion velocity of (2008MNRAS.389..113P; 2015MNRAS.449.1921P; 2015MNRAS.449.1954P; 2016MNRAS.456..853P; 2017MNRAS.471.4381S; 2020ApJ...889..170G; d2025arXiv250610700G). The chemical composition and the expansion velocity of the CSM lead to the interpretation that SNe Ibn arise from the terminal explosion of Wolf-Rayet (WR) stars (2007ApJ...657L.105F; 2008MNRAS.389..113P; 2017ApJ...836..158H; 2022A&A...658A.130D; 2022ApJ...927...25M). In particular, the pre-explosion outburst of SN 2006jc provided strong evidence for transitional objects between luminous blue variables and WR stars as the progenitors of SNe Ibn (2007Natur.447..829P). However, several issues remain unresolved to establish the connection between SNe Ibn and their possible progenitor WR stars. For instance, there is a discrepancy between the observed rate of SNe Ibn and the estimated explosion rate of SNe Ibn based on the Galactic WR stars, as well as between the inferred mass-loss rate of the CSM around SNe Ibn and WR stars (2007ARA&A..45..177C; 2014ARA&A..52..487S).
A more exotic category of interacting SNe, denoted as “Icn,” has recently been identified. The discovery event, SN 2019hgp, exhibits distinctive narrow emission lines of carbon and oxygen but lacks signatures of H and He (2022Natur.601..201G). Based on a detailed investigation of its circumstellar environment, 2022Natur.601..201G suggested that SN 2019hgp arises from the explosion of a WR star with a helium-poor and carbon-rich stellar wind. Up to now, only four events have been reported as SNe Icn (SNe 2019jc, 2021ckj, 2021csp, 2022ann; 2021arXiv210807278F; 2022ApJ...927..180P; 2022ApJ...938...73P; 2023MNRAS.523.2530D; 2023A&A...673A..27N). Despite the scarcity of high-quality data sets, SNe Icn present diverse properties in their light curves and host-galaxy environments, indicating multiple progenitor channels (2022ApJ...938...73P). In particular, SN 2019jc exhibits a relatively low peak luminosity and explosion energy (2022ApJ...938...73P), which may indicate alternative progenitor channels, such as the merger of an oxygen-neon white dwarf and a carbon-oxygen white dwarf (2024ApJ...967L..45W).
SN 2024abvb was discovered on UTC Nov. 22.33, 2024 (MJD 60636.334 2024TNSTR4579....1T) by the Asteroid Terrestrial-impact Last Alert System (ATLAS; 2018PASP..130f4505T). Follow-up spectroscopy obtained on MJD 60641.897 revealed a series of photoionized narrow C II emission lines superposed on a blue continuum, thereby putting it into the rare class of SNe Icn (2024TNSCR4674....1S). In Sec. II, we describe our multiband photometric and spectroscopic observations of SN 2024abvb and the data reduction. Sec. III presents a more detailed analysis of the light curve and spectra. Our discussion and conclusions are given in Sec. IV.
II Observations
We measure a radial velocity of km s-1 of SN 2024abvb by matching the centroids of a series of narrow C II emission lines (e.g., C II5890, C II 6580, and C II7235) in the classification spectrum (2024TNSCR4674....1S). We interpret the radial velocity as exclusively due to a redshift and adopt a Hubble constant of H km s-1 Mpc-1. We therefore derive a distance to SN 2024abvb of 165 Mpc assuming a standard CDM cosmology.
Adopting the position reported by the ATLAS discovery stream (RA , Dec. ; J2000) and assuming an extinction law as the Galactic average (1989ApJ...345..245C), we estimate a Galactic reddening of along the line of sight to SN 2024abvb by querying the NASA/IPAC NED Galactic Extinction calculator111https://ned.ipac.caltech.edu/forms/calculator.html and the extinction map by 2011ApJ...737..103S. We identify no obvious Na I D 5890, 5896 absorption doublet at the inferred redshift of the SN, whose line strength has been proposed as a tracer of the amount of the interstellar extinction (1997A&A...318..269M). As shown in Figure 1, we note that SN 2024abvb exploded in the outskirts of a galaxy whose centroid is measured as RA , Dec. (J2000), which yields a rather large separation between the SN and its tentative host (; and 23 kpc). We suggest that the line-of-sight reddening toward SN 2024abvb is dominated by the Galactic component and made no attempt to derive the extinction from its host.
II.1 Photometry
We present multiband light curves of SN 2024abvb in Figure 2 and list all the photometry in Table 1.
II.1.1 TNOT
The -band light curves of SN 2024abvb were obtained with the Tsinghua-Nanshan Optical Telescope (TNOT), an 80 cm reflector that began operating in 2024. TNOT is an equatorial telescope located at Nanshan Station of Xinjiang Astronomy Observatories, Chinese Academy of Sciences. The telescope system features an ASA800 Ritchey-Chrétien optical tube (/6.85, 5480 mm focal length) on ASA’s premium DDM500 equatorial mount. The field of view is designed to be . The Bessel system and the Sloan system are chosen as the primary filter set. The imaging system utilizes an Andor iKon-L DZ936 CCD camera with a pixel array, and the plate scale is 078 pixel-1.
All images obtained by TNOT were preprocessed using standard IRAF 222IRAF is distributed by the National Optical Astronomy Observatories, which are operated by the Association of Universities for Research in Astronomy, Inc., under cooperative agreement with the U.S. National Science Foundation (NSF). routines (1986SPIE..627..733T; 1993ASPC...52..173T), including bias subtraction and flat-field correction. For each epoch of the TNOT observation, 3–5 exposures were taken per filter. After computing the World Coordinate System solutions, we employed the reproject package333https://reproject.readthedocs.io/en/stable/index.html to align and combine the individual images. Point-spread-function (PSF) photometry was then performed using AutoPhOT (brennanAUTOmatedPhotometryTransients2022). Instrumental magnitudes of the SN were extracted using the AutoPhOT (brennanAUTOmatedPhotometryTransients2022) pipeline. In particular, for each frame, the PSF was constructed from bright, isolated field stars and matched to the SN. The fitting radius of the PSF model was set to the full width at half-maximum intensity (FWHM), and the background was estimated and subtracted iteratively during the PSF fit. For photometric calibration, we adopted the Pan-STARRS catalog for the , , and bands, and the APASS catalog for the Johnson and bands. Because the SN is bright and in the outskirts of its host galaxy, background removal by means of subtracting archival template images is not necessary.
II.1.2 ATLAS
We also include the - and -band data of SN 2024abvb from ATLAS using forced photometry444https://fallingstar-data.com/forcedphot/ (2018PASP..130f4505T; 2020PASP..132h5002S). Upper limits were derived based on a threshold above the flux uncertainties. Since the ALTAS footprint covered the field of SN 2024abvb –4 times per night, we stacked the nightly photometry to improve the signal-to-noise ratio (S/N).
II.1.3 Swift
The UltraViolet/Optical Telescope (UVOT) onboard the Neil Gehrels Swift Observatory conducted the first set of observations of SN 2024abvb on UTC 2024 Nov. 27 at 20:22:27.50 (MJD = 60641.85) in the , , , and optical bands (ObsID 00018944)(2004ApJ...611.1005G; 2005SSRv..120...95R), spanning from to days corresponding to the -band peak brightness. We performed aperture photometry using a Python-based pipeline that wraps around HEAsoft tools such as uvotmaghist and uvotimsum (2009AJ....137.4517B; 2014Ap&SS.354...89B).555The home page of the code on GitHub: https://github.com/gterreran/Swift_host_subtraction A 5′′-radius aperture on the target and a 20′′-radius aperture on the background are adopted for the measurements. Host-galaxy contamination at the SN location is minimal, so host subtraction was not applied to the photometry.
II.1.4 REM
The -band light curves of SN 2024abvb were obtained with the Rapid Eye Mount (REM) telescope (Covino:2004SPIE.5492.1613C; Zerbi:2004SPIE.5492.1590Z), which is a fully automated, 60 cm fast-response Ritchey-Chrétien’ telescope operating at the ESO La Silla Observatory in Chile. REM is equipped with two instruments: REMIR, a near-infrared (NIR) imaging camera (Conconi:2004SPIE.5492.1602C), and ROSS2, a visible-light camera (Tosti:2004SPIE.5492..689T). Both cameras simultaneously cover nearly the same field of view. With its dichroic beam-splitting design, REM can capture five images at once. Astrometric calibration of the optical and NIR images was performed using Astrometry.net (Lang:2010AJ....139.1782L). For each band, if more than two images with identical exposure times were available, they were combined using SWarp (swarp:2010ascl.soft10068B) to produce a deeper stacked image. Aperture photometry was then carried out on all images using SExtractor (sextracotr:1996A&AS..117..393B). Owing to the small field of view, photometric calibration was performed using an isolated bright reference star from the ATLAS-REFCAT2 catalog (Tonry:2018ApJ...867..105T).
II.2 Spectroscopy
The spectral time sequence of SN 2024abvb was obtained by the Xinglong 2.16 m telescope (XLT; 2016PASP..128k5005F), the Lijiang 2.4 m telescope (LJT; 2019RAA....19..149W), and the Shane 3 m telescope. Spectral reductions were performed using standard procedures within the IRAF software package, including steps such as bias subtraction, flat-field normalization, and cosmic-ray removal. Wavelength calibration was carried out using comparison-lamp exposures. Flux calibration was implemented using nightly exposures of spectrophotometric standard stars at air masses similar to those of SN 2024abvb. Telluric lines were removed by using an average telluric spectrum constructed from multiple observations of telluric standard stars scaled to match the flux level of the SN spectra. For comparison, we also present the classification spectrum obtained by the Nordic Optical Telescope (NOT) (2024TNSCR4674....1S).
The complete spectral sequence of SN 2024abvb discussed in this paper is shown in Figure 3. A log of spectroscopic observations is presented in Table 2.
III Analysis
III.1 Photometric Properties
As shown in Figure 2, we smoothed the multiband light curves of SN 2024abvb with a Gaussian process and estimated that the peak of the -band light curve is at MJD . Throughout the paper, all phases are given relative to the time of the -band peak brightness at MJD 60644.75. The last nondetection limit and the first detection in the ATLAS band yield mag at day 10.4 and mag at day 7.9, respectively. The former has been assigned to constrain the explosion time when fitting the multiband early light curves of SN 2024abvb (see Sec. III.3).
The -band light curve of SN 2024abvb reached a maximum of mag at MJD 60644.75, corresponding to an absolute magnitude of mag with the Galactic extinction correction. The -band peak brightness of SN 2024abvb is comparable to that of SN 2021csp (2022ApJ...927..180P), both of which represent cases at the luminous end of the SN Icn distribution, as illustrated in the upper panel of Figure 4. Notably, SN 2024abvb exhibits a rapid drop after days compared to SN 2021csp. Except for an overall mag excess in brightness, the light-curve morphology of SN 2024abvb more resembles that of SN 2021ckj. The latter overlaps with the template light curve of SNe Ibn as presented by the gray-shaded band in the upper panel of Figure 4.
In the bottom panel of Figure 4 we present the color evolution of SN 2024abvb spanning days before to 37 days after the -band light-curve peak. A monotonic change in color from to 0.0 mag can be identified in the time interval to 16 days during which the multiband light curves are well sampled.
III.2 Spectroscopic Properties
The early-time spectra of SN 2024abvb are characterized by three emission features at rest-frame wavelengths near 5890 Å, 6580 Å, and 7235 Å superimposed on a blue continuum as shown in Figure 3. Portraits of photoionized spectral features centered at several C II features, whose rest-frame wavelengths are 5890 Å, 6578 Å, and 7234 Å, are shown respectively in the left, middle, and right columns of Figure 5. These photonionization-powered narrow emission lines, whose FWHM are found to be km s-1, have vanished by day . Limited by the insufficient S/N, we do not attempt to investigate the temporal evolution of the width of these short-lived emission lines. Despite the proximity to He I 5876, owing to the lack of other features of He I such as near 6678 and 7065 Å, we attribute the emission profile presented in the left column of Figure 5 to C II 5890 and conclude the absence of helium in the early-phase spectra of SN 2024abvb.
Figure 6 compares the spectra of SN 2024abvb at days 1.2 and 3.7 with the spectra of a selected sample of well-observed SNe Ibn/Icn at similar phases. The spectral features of SN 2024abvb exhibit a strong resemblance to those of SNe Icn, in which the C II emission lines are prominent and helium lines are absent. We also note that, unlike other SNe Icn, the early-time spectra of SN 2024abvb display no sign of any C III feature. A natural explanation for this discrepancy could be a relatively lower temperature in the photoionized CSM of SN 2024abvb, where the highest temperature produces the highest excitation states.
III.3 Modeling with a Hybrid Model
Multiple energy sources may contribute to the early emission of core-collapse SNe. These include the prompt release of electromagnetic emission ahead of the shock wave that ejects the stellar envelope (namely the shock-breakout emission), the recombination emission that mostly takes place in the outermost H-rich envelope and manifests as strong Balmer lines, the decay chain of explosion-synthesized radioactive nuclei (in particular 56NiCoFe), the conversion of kinetic energy to radiation facilitated by the SN ejecta expanding into the CSM, and the injection of energy from a central engine such as a magnetar as it spins down (e.g., 1980ApJ...237..541A; 1982ApJ...253..785A; 1994ApJ...420..268C; 2010ApJ...717..245K; 2010ApJ...725..904N; 2011ApJ...729...61B; 2012ApJ...746..121C; 2013MNRAS.435.1520M; 2016A&A...589A..53N). For SN 2024abvb, all these aforementioned mechanisms, except H recombination, remain viable candidates. Detecting prominent emission-line features in SN 2024abvb provides direct observational evidence of CSM interaction. The lack of early-time observations hinders a comprehensive modeling of the shock-cooling process. Furthermore, we did not attempt to model the luminosity evolution by introducing any central engine. Therefore, our final hybrid model incorporates the CSM interaction and the radioactive decay of 56Ni, both implemented through the Modular Supernova Fitting Tool (MOSFiT; 2018ApJS..236....6G). The best-fit parameters and the associated 1 uncertainties were derived using a Markov Chain Monte Carlo (MCMC) approach.
Within the MOSFiT framework, the 56Ni decay component is parameterized by the ejecta mass (), the kinetic energy (), and the 56Ni mass (). Meanwhile, the ejecta-CSM interaction is characterized by , , and the mass of the CSM (). The MCMC fitting procedure samples a uniform prior distribution for all fitting parameters within the following ranges: (0.01, 30) , (0.1, 30) , (0.05, 5) , (0.0001, 1.0) , and (, ) days. In a parallel configuration, we set walk to 150 and iteration to 5000. The MCMC posterior distributions of the derived parameters and the fit to the multiband light curves of SN 2024abvb are presented in Figures 7 and 8, respectively. In particular, the best fit to the 30 day light curves gives masses of the ejecta, the CSM, and the radioactive nickel of , , and solar masses, respectively. The derived parameters are consistent with those estimated from other well-studied cases (e.g., 2022ApJ...927..180P; 2022ApJ...938...73P; 2023MNRAS.523.2530D).
The posterior distribution of as presented in the bottom row of Figure 7 shows a long tail toward higher masses. The rather poor constraint on the 56Ni mass is due to the lack of photometry after day, when the relative energy contribution by the 56CoFe progressively increases. As illustrated in the lower-right panel of Figure 8, the radiation around peak brightness is mainly produced by the ejecta-CSM interaction compared to that by the 56Ni decay. As a result, the near-peak photometry itself cannot place stringent constraints on the mass of radioactive nickel synthesized in the ejecta of the strongly interacting SN 2024abvb.
In the model adapted to the MOSFiT framework, we assume a spherically symmetric emitting zone for the radioactive decay of nickel, and the radiation diffusion throughout the SN ejecta is described following the prescriptions of Eq. 9 of 2012ApJ...746..121C. We do not consider the diffusion process within the unshocked CSM for the emission generated by either the 56Ni decay or the ejecta-CSM interaction. Such a simplified approach may introduce systematic uncertainties in estimating the explosion parameters, for example, the explosion time (2025ApJ...984...44H). The exact process by which a photon propagates through the SN ejecta and diffuses out of the surrounding CSM requires a more careful treatment of model parameters, such as the density profile, the time-varying optical depth, and the geometry of both the ejecta and the CSM. The calculation also demands computationally intensive Monte Carlo radiative-transfer simulations. Detailed modeling of the photometric evolution of SN 2024abvb is beyond the scope of this paper.
We also note that our model assumes a homogeneous mass-loss wind from the progenitor star. This isotropic CSM content was found in contradiction to the aspherical CSM directly adjacent to the progenitor from polarimetry (e.g., 2023MNRAS.519.1618Y; 2023ApJ...955L..37V; 2025arXiv250503975V). In some cases, the radial density structure of the CSM deviates from a steady, spherically symmetric mass-loss profile, indicating enhanced episodes of mass loss or a series of mass ejections during outbursts of luminous blue variables (e.g., 1994A&A...290..819L; 2008A&A...483L..47T). Therefore, we consider the estimated CSM mass a plausible upper limit.
IV Discussion and Conclusions
In this paper, we present photometry and spectroscopy of SN 2024abvb, which we classify as an SN Icn based on the identification of a set of relatively narrow (FWHM –2000 km s-1), singly ionized carbon emission lines. Such a carbon-rich CSM expanding at a fairly high velocity implies a hydrogen- and helium-depleted, carbon-dominant, Wolf-Rayet-star-like wind (WC type). Among the SN Icn subclass, the light-curve morphology exhibits significant diversity with an -band peak magnitude ranging from 17 mag (SN 2019jc, 2022ApJ...938...73P) to 20 mag (SN 2021csp, 2022ApJ...927..180P; 2022ApJ...938...73P). SN 2024abvb shows a peak brightness similar to that of SN 2021csp but declines faster.
We fit the UV-to-optical light curves of SN 2024abvb, adopting a hybrid analytic model that accounts for both the ejecta-CSM interaction and the radioactive decay of 56Ni. The best-fit result suggests a rather small nickle mass of M⊙ in the ejecta, and the near-peak emission of SN 2024abvb is dominated by the ejecta-CSM interaction. This low Ni mass is consistent with the analysis of other SNe Icn.
The nature of SNe Icn still remains controversial. Similar to SNe Ibn, they may originate from the explosion of WR stars, since both exhibit signatures of CSM that expand at rather high velocities of and have stripped their outer H/He envelopes. The mechanism driving such intense ejection of the outer H/He envelopes immediately prior to the terminal explosion remains unexplained. The absence of apparent P Cygni profiles in our early-time spectral sequence of SN 2024abvb hinders the measurement of the wind velocity. The modeled () and () are smaller than those of other SNe Icn or stripped-envelope SNe as shown in Figure 9, indicating that the progenitor star of SN 2024abvb experienced significant mass loss prior to the explosion.
SN 2024abvb exploded in the outskirts of its host galaxy, which may correspond to a relatively long evolving history for its progenitor system. One plausible explanation for the small ejecta mass and nickel mass associated with the bright, fast-declining light curve of SN 2024abvb is that the exploding progenitor star originated in a binary system consisting of a helium star and a compact companion (2015MNRAS.451.2123T). The H and He envelopes of the former are progressively stripped by the latter through Roche-lobe overflow (2007ASPC..372..397M). Such enhanced mass loss toward the orbital plane of the binary system would take place when the Roche lobe is filled, and is suggested to be capable of efficiently stripping the H envelope (2014ARA&A..52..487S).
The chemical composition and wind velocity of the CSM encode information about the mass-loss history of massive stars nearing the end of their lives. The fusion reactions transform massive stars into a structure of concentric shells. One of the key deliverables of stripped-envelope SNe is testing the long-hypothesized internal layering that develops at the end of the evolution of massive stars: progressively heavier elements are synthesized toward the iron core at the center. Spectrophotometric observations from the earliest possible phases would reveal the composition and physical conditions of the surface layers of the exploding star, which are immediately expelled before the explosion. Observations of Type Ibn/Icn SNe, together with the very recently reported Type Ien SN 2021yfj (2024TNSAN.239....1G; 2024TNSAN.240....1G; 2024arXiv240902054S), whose immediate CSM exhibits chemical signatures of silicon, sulfur, and argon, provide key clues for understanding the nature and process of such a stripping sequence emerging from the death of massive stars.
| MJD | Phase | Filter | Magnitude | Instrument | System |
| 60642.601 | -2.149 | 17.100.10 | TNOT | Vega | |
| 60642.602 | -2.148 | 17.100.08 | TNOT | Vega | |
| 60642.604 | -2.146 | 17.070.06 | TNOT | Vega | |
| 60642.605 | -2.145 | 17.090.08 | TNOT | Vega | |
| 60643.577 | -1.174 | 17.030.10 | TNOT | Vega | |
| 60643.578 | -1.172 | 17.030.13 | TNOT | Vega | |
| 60643.579 | -1.171 | 17.050.06 | TNOT | Vega | |
| 60643.580 | -1.170 | 17.050.11 | TNOT | Vega | |
| 60644.673 | -0.077 | 17.020.13 | TNOT | Vega | |
| 60644.675 | -0.075 | 17.010.08 | TNOT | Vega | |
| 60644.676 | -0.074 | 17.020.07 | TNOT | Vega | |
| 60644.677 | -0.073 | 17.020.07 | TNOT | Vega | |
| 60649.611 | 4.861 | 17.350.18 | TNOT | Vega | |
| 60650.646 | 5.896 | 17.270.14 | TNOT | Vega | |
| 60651.557 | 6.808 | 17.490.08 | TNOT | Vega | |
| 60652.599 | 7.849 | 17.610.12 | TNOT | Vega | |
| No. | UTC Date | MJD | Phase (d) | Exp. (s) | Telescope+Inst. | Range (Å) | Airmass |
| 1 | 11.27 2024 | 60641.90 | 1800 | NOT+ALFOSC | 3700-8000 | / | |
| 2 | 11.29 2024 | 60643.59 | 1800 | LJT+YFOSC | 3600-8000 | 1.2 | |
| 3 | 11.30 2024 | 60644.61 | 3000 | XLT+BFOSC | 4000-8000 | 2.1 | |
| 4 | 12.01 2024 | 60645.65 | 0.9 | 1800 | LJT+YFOSC | 3600-8000 | 1.4 |
| 5 | 12.02 2024 | 60646.51 | 1.8 | 3600 | XLT+BFOSC | 4000-8000 | 1.5 |
| 6 | 12.03 2024 | 60647.66 | 2.9 | 3600 | XLT+BFOSC | 4000-8000 | 3.6 |
| 7 | 12.04 2024 | 60648.49 | 3.7 | 3600 | XLT+BFOSC | 3750-8000 | 1.4 |
| 8 | 12.05 2024 | 60649.68 | 4.9 | 1700 | LJT+YFOSC | 3600-8000 | 1.7 |
| 9 | 12.08 2024 | 60652.68 | 7.9 | 1500 | LJT+YFOSC | 3600-8000 | 1.8 |
| 10 | 12.10 2024 | 60654.23 | 9.5 | 3600 | Shane+Kast | 4000-8000 | 1.5 |
| 11 | 12.18 2024 | 60662.55 | 17.8 | 2400 | LJT+YFOSC | 4000-8000 | 1.2 |